# Can Graph Learning Improve Planning in LLM-based Agents?

### v2

@@ 2.2 Current LLM-based Solution to Task Planning
Figure 1: Illustration of Task Planning in Language Agents (e.g., HuggingGPT)
@@ 4.2 A Training-free GNN-based Approach
For each decomposed step outputted by the first stage, we use a GNN to select a corresponding node within the task graph. Suppose we are selecting the node for the $i$-th decomposed step. First, we utilize a small pre-trained language model, e5-355M [61], to embed the $i$-th decomposed step. The resulting embedding is denoted as $\bm{x}_{i}^{\text{step}}$. Second, for the task graph, we first use the same pre-trained language model e5-355M to convert each node’s description into embeddings, denoted as the node feature $\bm{h}^{0}_{v}$, where the superscript indicates the layer and the subscript represents the node. Then we adopt a $K$-layer SGC [64] to compute the final node embeddings, resulting in $\bm{h}_{v}=\bm{h}_{v}^{(K)}$. Given a sequence of previously selected task nodes $\{v_{1},\cdots,v_{i-1}\}$, the next node $v_{i}$ is chosen according to $v_{i}=\text{argmax}_{v\in\mathcal{N}(v_{i-1})}\;\langle\bm{h}_{v},\bm{x}_{i}^{\text{step}}\rangle$, where $\bm{h}_{v}$ is the final node embedding, and $\mathcal{N}(v_{i-1})$ denotes the neighbors of node $v_{i-1}$ in the task graph. Particularly, $v_{1}$ can be selected from the whole graph.
As e5-355M is pre-trained and SGC is parameter-free, the proposed method requires no additional training and can be effectively applied in a zero-shot manner.
@@ 4.3 A Training-required GNN-based Approach
Training Loss: The problem in the dataset can be viewed as a binary ranking problem, where the labeled node is $1$ and the other nodes are 00. Therefore, we adopt the Bayesian Personalized Ranking (BPR) loss [40] designed for recommendation with binary rankings. The loss function is given by $\ell=\sum_{(\bm{x}^{\text{step}},v,v^{\prime})}-\log\;\sigma(\langle\bm{h}_{v},\bm{x}^{\text{step}}\rangle-\langle\bm{h}_{v^{\prime}},\bm{x}^{\text{step}}\rangle)$, where $\bm{x}^{\text{step}}$ represents the embedding of the step’s textual description generated by e5-355M, $v$ is the ground-truth task, and $v^{\prime}$ is a negative task. We select negative tasks that are textually similar to the positive task, and for computational efficiency, we limit our selection to $2$ negative tasks per positive task. The trainable parameters may merely include GNNs or both GNNs and e5-355M with illustrative configurations shown in Figure 8 in Appendix.
@@ 5.2 Performance of the Training-free Approach
We compare the performance across three training-free methods: (1) LLM’s Direct Inference is introduced in Section 2.2. (2) GraphSearch [32, 51, 31] leverages the classic graph search method to generate the candidate nodes and use LLMs to give a score for node selection. Given a step, GreedySearch consistently selects the node with the highest score and adjacent to the previous task node; AdaptiveSearch selects the nodes with scores above a fixed threshold, adjusting the breadth of the search space in an adaptive mode; BeamSearch retains the $k$ nodes with highest scores. (3) SGC [64] employs a training-free SGC for task retrieval based on decomposed task steps. The details of baselines are given in Section E.1 and illustrated in Figure 6. Table 1 shows both the overall performance and token consumption costs, with results of Accuracy (Acc) moved to Table 9 in Appendix.
@@
Compared with direct inference, integrating an SGC consistently improves performance, underscoring the effectiveness of the proposed method. GraphSearch-type methods rely on beam search to identify paths and employ LLMs for evaluation, where longer processing times generally lead to better outcomes. Notably, our proposed method achieves comparable or superior performance (Table 1 and Table 9) to BeamSearch while requiring $5$-$10$ times fewer tokens (Table 1) and inference time (Table 10). The case studies are provided in Appendix H. However, we observed only marginal improvements with GPT-4-turbo. A unique feature of GPT-4-turbo is its ability to manage ChatGPT-plugins, and it may have been specially trained on task planning datasets. In addition, the pre-trained language model used for feature extraction in SGC is e5-355M, which may not be sufficiently powerful to effectively analyze GPT-4’s output. A detailed diagnostic analysis of cases involving GPT-4 is provided in Figure 11 in Appendix.
@@ 5.3 Performance of the Training-based Approaches
Settings: We further explore the efficacy of training-based GNNs in three TaskBench datasets. The TMDB dataset is excluded due to its limited sample size. Throughout our experiments, we trained a spectrum of GNN variants, both with and without co-training the small LM (i.e., e5-355M), whose role is to generate node embeddings derived from task names and descriptions. Owing to space constraints, we only show the performance of GraphSAGE in the main text, relegating a detailed comparison of all the situations to Table 11 and Table 12 in Appendix.
@@
Observations: From Table 2, we observe a significant improvement in performance when employing a training-based GraphSAGE approach over the training-free method. However, the co-training of GNNs with e5-355M does not yield a marked improvement, suggesting that message passing is the crucial element for enhancing performance. Further analysis across a broad spectrum of GNNs (as shown in Table 11 and Table 12) reveals that powerful GNNs, such as GINs, perform similarly to networks perceived as less complex, like GCNs, and even underperform compared to GraphSAGE. This pattern indicates that the task’s challenge may not lie in the expressiveness of the models but rather in their ability to generalize. Regarding baselines, TAPE is unsuitable for task planning as its classification approach simplifies task planning, overlooking task dependencies. While GraphToken demonstrates superior performance over LLMs’ direct inference, we have noted instances of minor hallucination. This observation suggests that GraphToken’s understanding of the task graph is not yet perfect. In addition, GraphToken is limited to open-sourced LLMs. The proposed approaches also greatly improve the parameter prediction performance, as given in Appendix G (e.g. $9\%$ improvement to to GPT-3.5 and $3\%$ improvement to GPT-4).
Efficiency: The details of the training time are given in Table 13. The training cost is remarkably low because we use e5-355M [61] as the text embedding model for GNNs. If the trainable parameters are limited to the GNNs alone, training typically concludes within just 3 minutes. Furthermore, the training duration extends to only 15 minutes when GNNs are jointly trained with e5-355M model. This efficiency stands in stark contrast to the 10-20 hours required for tuning open-sourced LLMs.
@@ A.4 LLMs for Graphs
Considering these negative results, a new line of research has emerged that utilizes the output of GNNs as tokens for LLMs, as seen in GraphGPT [52], GraphLLM [9], and GraphToken [37]. These approaches have demonstrated significant improvements in performance on GNN-related tasks. However, they have not yet been applied to task planning due to the lack of extensive training data. A promising future direction involves using task planning data generated by GPT to fine-tune graph foundation models, such as GraphGPT [52], and applying them to task planning. This paper proposes to use task planning as a new benchmark for this line of research.
@@ E.1 Implementation of Baselines
Ours SGC: Regarding the choices of LM backbones, for integrating GPT-3.5-turbo and GPT-4-turbo with SGC, the Roberta-355M [39] serves as the text encoder. For all other datasets and LLMs, the e5-355M [61] configuration is employed.
@@ F.1 Implementation of Training-based GNNs
LM and GNN Configuration: For training-based GNNs, we uniformly use the e5-355M model [61] as the LM backbone. For the graph encoder, our setup includes a single layer with a hidden dimension of $1024$. During the model training, we set the batch size to $512$ and run for $20$ epochs with a learning rate of $31e-$ 3. We use the Adam optimizer [24] and implement an early stopping mechanism with a patience of $5$ epochs to prevent over-fitting. All experiments are conducted on a single NVIDIA A100-80G GPU.
@@ F.2 Implementation of TAPE and GraphToken
TAPE: To adapt TAPE for task planning, we reformulate the planning task as a node classification problem, aiming to classify a user request into the appropriate task labels within the task graph. Firstly, LLMs interpret user requests by generating high-quality explanatory text, i.e., chain-of-thought reasoning, to understand each request. Then, a fine-tuned LM encodes these chain-of-thought texts into latent embeddings. The fine-tuning process is based on pairs of textual descriptions and their corresponding ground-truth tasks. Finally, GNNs select the suitable tasks by leveraging both the generated text embeddings and the task embeddings. For a fair performance comparison, we fine-tune the LM as e5-355M and configure the GNN as a $2$-layer GraphSAGE with a hidden dimension of $1024$. The training data for both the LM and GNNs are consistent, utilizing the same split of the training dataset.
GraphToken: For adapting GraphToken, we first encode the task graph $G(V,E,X)$ by computing each node’s representation $\bm{x}_{v}$ by feeding its descriptive text into the pre-trained LM, e5-355M. Then, a GNN transforms the initial node embedding matrix into $H$. The GNN is configured as a $2$-layer GraphSAGE with a hidden dimension of $1024$. A mean pooling operation is further applied to the node embeddings to obtain the graph’s overall representation as $\bm{h}_{G}=\text{Mean-Pooling}(H)$. Finally, both $\bm{h}_{G}$ and the text embeddings of input instruction and user request are concatenated and fed into the LLM to generate the output planning result. During this process, the backbone LLMs are frozen, and only the GNN’s parameters are tuned. For each dataset, we train over $4$ epochs using the same $3,000$ training samples consisting of <user request, ground-truth planning> pairs. The batch size is set to $16$, with a maximum input length of $512$ and a maximum output length of $300$ for HuggingFace and Multimedia datasets, and $600$ for DailyLife dataset. The learning rate is configured to $51e-$ 5.
@@ F.6 Computational Cost Analysis
Efficiency of Training GNNs: During experiments, each dataset shares the same GNN configuration: $1$ single layer with a hidden dimension of $1024$. Therefore, for each GNN, its number of parameters remains consistent across datasets. The parameter scales for GNN variants range from 1M to 4M, and the total training time for each dataset requires only 2-4 minutes, comparable to the time taken by GraphSearch to fulfill a single request. For LM+GNN co-trained mode, where e5-355M [61] serves as the LM backbone, training times increase to approximately 6-12 minutes. In summary, both modes demonstrate high efficiency, with total training times spanning minutes, showcasing their ability to rapidly adapt to new task planning scenarios.
@@ G.2 Empirical Results of LLMs Predicted Parameters
Observation: From the results shown in Table 14, we can conclude that: (1) LLMs’ direct inference of invocation arguments is unsatisfactory. Even for the strongest LLM, GPT-4-turbo, the Parameter-Value F1 score is only $55\%$, which is far from expectations. (2) With accurate planning provided by GNNs, LLMs can leverage their inherent reasoning abilities to analyze the context and correctly fill in the parameters. Across four LLMs, improvements of $2\%$ to $20\%$ can be observed, highlighting the advantages of GNN-enhanced planning. (3) Our method is well-suited for training-free scenarios, as the tasks retrieved by the training-free SGC already enable LLMs to infer the parameters effectively, with only a minor performance gap compared to the training-required GraphSAGE.

### v3

@@ 2.2 Current LLM-based Solution to Task Planning
Figure 1: Illustration of Task Planning in Language Agents (e.g., HuggingGPT [46])
@@ 4.2 A Training-free GNN-based Approach
For each decomposed step outputted by the first stage, we use a GNN to select a corresponding node within the task graph. Suppose we are selecting the node for the $i$-th decomposed step. First, we utilize a small pre-trained language model, e5-335M [62], to embed the $i$-th decomposed step. The resulting embedding is denoted as $\bm{x}_{i}^{\text{step}}$. Second, for the task graph, we first use the same pre-trained language model e5-335M to convert each node’s description into embeddings, denoted as the node feature $\bm{h}^{0}_{v}$, where the superscript indicates the layer and the subscript represents the node. Then we adopt a $K$-layer SGC [65] to compute the final node embeddings, resulting in $\bm{h}_{v}=\bm{h}_{v}^{(K)}$. Given a sequence of previously selected task nodes $\{v_{1},\cdots,v_{i-1}\}$, the next node $v_{i}$ is chosen according to $v_{i}=\text{argmax}_{v\in\mathcal{N}(v_{i-1})}\;\langle\bm{h}_{v},\bm{x}_{i}^{\text{step}}\rangle$, where $\bm{h}_{v}$ is the final node embedding, and $\mathcal{N}(v_{i-1})$ denotes the neighbors of node $v_{i-1}$ in the task graph. Particularly, $v_{1}$ can be selected from the whole graph.
As e5-335M is pre-trained and SGC is parameter-free, the proposed method requires no additional training and can be effectively applied in a zero-shot manner.
@@ 4.3 A Training-required GNN-based Approach
Training Loss: The problem in the dataset can be viewed as a binary ranking problem, where the labeled node is $1$ and the other nodes are 00. Therefore, we adopt the Bayesian Personalized Ranking (BPR) loss [41] designed for recommendation with binary rankings. The loss function is given by $\ell=\sum_{(\bm{x}^{\text{step}},v,v^{\prime})}-\log\;\sigma(\langle\bm{h}_{v},\bm{x}^{\text{step}}\rangle-\langle\bm{h}_{v^{\prime}},\bm{x}^{\text{step}}\rangle)$, where $\bm{x}^{\text{step}}$ represents the embedding of the step’s textual description generated by e5-335M, $v$ is the ground-truth task, and $v^{\prime}$ is a negative task. We select negative tasks that are textually similar to the positive task, and for computational efficiency, we limit our selection to $2$ negative tasks per positive task. The trainable parameters may merely include GNNs or both GNNs and e5-335M with illustrative configurations shown in Figure 8 in Appendix.
@@ 5.2 Performance of the Training-free Approach
We compare the performance across three training-free methods: (1) LLM’s Direct Inference is introduced in Section 2.2. (2) GraphSearch [33, 52, 32] leverages the classic graph search method to generate the candidate nodes and uses LLMs to give a score for node selection. Given a step, GreedySearch consistently selects the node with the highest score and adjacent to the previous task node; AdaptiveSearch selects the nodes with scores above a fixed threshold, adjusting the breadth of the search space in an adaptive mode; BeamSearch retains the $k$ nodes with highest scores. (3) SGC [65] employs a training-free SGC for task retrieval based on decomposed task steps. The details of baselines are given in Section E.1 and illustrated in Figure 6. Table 1 shows both the overall performance and token consumption costs, with results of Accuracy (Acc) moved to Table 9 in Appendix.
@@
Compared with direct inference, integrating an SGC consistently improves performance, underscoring the effectiveness of the proposed method. GraphSearch-type methods rely on beam search to identify paths and employ LLMs for evaluation, where longer processing times generally lead to better outcomes. Notably, our proposed method achieves comparable or superior performance (Table 1 and Table 9) to BeamSearch while requiring $5$-$10$ times fewer tokens (Table 1) and inference time (Table 10). The case studies are provided in Appendix H. However, we observed only marginal improvements with GPT-4-turbo. A unique feature of GPT-4-turbo is its ability to manage ChatGPT-plugins, and it may have been specially trained on task planning datasets. In addition, the pre-trained language model used for feature extraction in SGC is e5-335M, which may not be sufficiently powerful to effectively analyze GPT-4’s output. A detailed diagnostic analysis of cases involving GPT-4 is provided in Figure 11 in Appendix.
@@ 5.3 Performance of the Training-based Approaches
Settings: We further explore the efficacy of training-based GNNs in three TaskBench datasets. The TMDB dataset is excluded due to its limited sample size. Throughout our experiments, we trained a spectrum of GNN variants, both with and without co-training the small LM (i.e., e5-335M), whose role is to generate node embeddings derived from task names and descriptions. Owing to space constraints, we only show the performance of GraphSAGE in the main text, relegating a detailed comparison of all the situations to Table 11 and Table 12 in Appendix.
@@
Observations: From Table 2, we observe a significant improvement in performance when employing a training-based GraphSAGE approach over the training-free method. However, the co-training of GNNs with e5-335M does not yield a marked improvement, suggesting that message passing is the crucial element for enhancing performance. Further analysis across a broad spectrum of GNNs (as shown in Table 11 and Table 12) reveals that powerful GNNs, such as GINs, perform similarly to networks perceived as less complex, like GCNs, and even underperform compared to GraphSAGE. This pattern indicates that the task’s challenge may not lie in the expressiveness of the models but rather in their ability to generalize. Regarding baselines, TAPE is unsuitable for task planning as its classification approach simplifies task planning, overlooking task dependencies. While GraphToken demonstrates superior performance over LLMs’ direct inference, we have noted instances of minor hallucination. This observation suggests that GraphToken’s understanding of the task graph is not yet perfect. In addition, GraphToken is limited to open-sourced LLMs. Besides, our proposed approaches also greatly boost the parameter prediction performance, as given in Appendix G (e.g. $9\%$ improvement to GPT-3.5-turbo and $3\%$ improvement to GPT-4-turbo).
Efficiency: The details of the training time are given in Table 13. The training cost is remarkably low because we use e5-335M [62] as the text embedding model for GNNs. If the trainable parameters are limited to the GNNs alone, training typically concludes within just 3 minutes. Furthermore, the training duration extends to only 15 minutes when GNNs are jointly trained with e5-335M model. This efficiency stands in stark contrast to the 10-20 hours required for tuning open-sourced LLMs.
@@ A.4 LLMs for Graphs
Considering these negative results, a new line of research has emerged that utilizes the output of GNNs as tokens for LLMs, as seen in GraphGPT [53], GraphLLM [9], GraphToken [38], and G-Retriever [17]. These approaches have demonstrated significant improvements in performance on GNN-related tasks. However, they have not yet been applied to task planning due to the lack of extensive training data. A promising future direction involves using task planning data generated by GPT to fine-tune graph foundation models, such as GraphGPT [53], and applying them to task planning. This paper proposes to use task planning as a new benchmark for this line of research.
@@ E.1 Implementation of Baselines
Ours SGC: Regarding the choices of LM backbones, for integrating GPT-3.5-turbo and GPT-4-turbo with SGC, the Roberta-355M [40] serves as the text encoder. For all other datasets and LLMs, the e5-335M [62] configuration is employed.
@@ F.1 Implementation of Training-based GNNs
LM and GNN Configuration: For training-based GNNs, we uniformly use the e5-335M model [62] as the LM backbone. For the graph encoder, our setup includes a single layer with a hidden dimension of $1024$. During the model training, we set the batch size to $512$ and run for $20$ epochs with a learning rate of $31e-$ 3. We use the Adam optimizer [25] and implement an early stopping mechanism with a patience of $5$ epochs to prevent over-fitting. All experiments are conducted on a single NVIDIA A100-80G GPU.
@@ F.2 Implementation of TAPE and GraphToken
TAPE: To adapt TAPE for task planning, we reformulate the planning task as a node classification problem, aiming to classify a user request into the appropriate task labels within the task graph. Firstly, LLMs interpret user requests by generating high-quality explanatory text, i.e., chain-of-thought reasoning, to understand each request. Then, a fine-tuned LM encodes these chain-of-thought texts into latent embeddings. The fine-tuning process is based on pairs of textual descriptions and their corresponding ground-truth tasks. Finally, GNNs select the suitable tasks by leveraging both the generated text embeddings and the task embeddings. For a fair performance comparison, we fine-tune the LM as e5-335M and configure the GNN as a $2$-layer GraphSAGE with a hidden dimension of $1024$. The training data for both the LM and GNNs are consistent, utilizing the same split of the training dataset.
GraphToken: For adapting GraphToken, we first encode the task graph $G(V,E,X)$ by computing each node’s representation $\bm{x}_{v}$ by feeding its descriptive text into the pre-trained LM, e5-335M. Then, a GNN transforms the initial node embedding matrix into $H$. The GNN is configured as a $2$-layer GraphSAGE with a hidden dimension of $1024$. A mean pooling operation is further applied to the node embeddings to obtain the graph’s overall representation as $\bm{h}_{G}=\text{Mean-Pooling}(H)$. Finally, both $\bm{h}_{G}$ and the text embeddings of input instruction and user request are concatenated and fed into the LLM to generate the output planning result. During this process, the backbone LLMs are frozen, and only the GNN’s parameters are tuned. For each dataset, we train over $4$ epochs using the same $3,000$ training samples consisting of <user request, ground-truth planning> pairs. The batch size is set to $16$, with a maximum input length of $512$ and a maximum output length of $300$ for HuggingFace and Multimedia datasets, and $600$ for DailyLife dataset. The learning rate is configured to $51e-$ 5.
@@ F.6 Computational Cost Analysis
Efficiency of Training GNNs: During experiments, each dataset shares the same GNN configuration: $1$ single layer with a hidden dimension of $1024$. Therefore, for each GNN, its number of parameters remains consistent across datasets. The parameter scales for GNN variants range from 1M to 4M, and the total training time for each dataset requires only 2-4 minutes, comparable to the time taken by GraphSearch to fulfill a single request. For LM+GNN co-trained mode, where e5-335M [62] serves as the LM backbone, training times increase to approximately 6-12 minutes. In summary, both modes demonstrate high efficiency, with total training times spanning minutes, showcasing their ability to rapidly adapt to new task planning scenarios.
@@ G.2 Empirical Results of LLMs Predicted Parameters
Observation: From the results shown in Table 14, we can conclude that: (1) LLMs’ direct inference of invocation arguments is unsatisfactory. Even for the strongest LLM, GPT-4-turbo, the Parameter-Value F1 score is only $55\%$, which is far from expectations. (2) With accurate planning provided by GNNs, LLMs can leverage their inherent reasoning abilities to analyze the context and correctly fill in the parameters. Across four LLMs, improvements of $3\%$ to $23\%$ can be observed, highlighting the advantages of GNN-enhanced planning. (3) Our method is well-suited for training-free scenarios, as the tasks retrieved by the training-free SGC already enable LLMs to infer the parameters effectively, with only a minor performance gap compared to the training-required GraphSAGE.